OpenStudy (anonymous):

Need help evaluating this integral.. I know the answer, just not how to get to it

OpenStudy (anonymous):

\[\int\limits_{?}^{?}x^2 /(x^4-1)dx\]

myininaya (myininaya):

did you try integration by parts?

OpenStudy (anonymous):

Yes that did not help..

myininaya (myininaya):

how far did you get? did you get this far: \[\frac{x^2}{x^4-1}=\frac{\frac{1}{4}}{x-1}+\frac{\frac{-1}{4}}{x+1}+\frac{\frac{1}{2}}{x^2+1} ?\]

myininaya (myininaya):

Or is that the problem getting there?

OpenStudy (anonymous):

yeah the problem is getting there..

myininaya (myininaya):

\[x^4-1=(x^2-1)(x^2+1)=(x-1)(x+1)(x^2+1)\] First factor the bottom as much as you can over the rationails

myininaya (myininaya):

rationals *

OpenStudy (anonymous):

alright that makes sense so far

myininaya (myininaya):

the linear factors we will do constant/the linear factor the non linear we will do (Constant*x+constant)/the non linear factor

myininaya (myininaya):

\[\frac{x^2}{x^4-1}=\frac{A}{x-1}+\frac{B}{x+1}+\frac{Cx+D}{x^2+1}\]

myininaya (myininaya):

So these are equal expressions...

OpenStudy (anonymous):

Yeah the Cx+D was mostly throwing me off

myininaya (myininaya):

That means if we combine the fractions on the right hand side then the top should be the same as the top as the left and bottom will be the same as the bottom on the other side

myininaya (myininaya):

\[\frac{x^2}{x^4-1}=\frac{A(x+1)(x^2+1)+B(x-1)(x^2+1)+(Cx+D)(x-1)(x+1)}{(x-1)(x+1)(x^2+1)}\]

myininaya (myininaya):

\[\frac{x^2}{x^4-1}=\frac{A(x+1)(x^2+1)+B(x-1)(x^2+1)+(Cx+D)(x-1)(x+1)}{x^4-1} \] The bottoms are the same The tops should be the same but we need to find A,B,C,D such that it is

OpenStudy (anonymous):

Alright I think I got it now. So I get D to=1/2, 1/4 and b=-1/4 then i just put bcak into equation, take the integral and im done?

myininaya (myininaya):

\[1 \cdot x^2+A(x^3+x+x^2+1)+B(x^3+x-x^2-1)+(Cx+D)(x^2-1)\]

myininaya (myininaya):

what about A? C=0 by the way

myininaya (myininaya):

oh yes are you saying A is 1/4?

OpenStudy (anonymous):

yes

myininaya (myininaya):

right! :)

OpenStudy (anonymous):

While your here do you mind telling me how i did this one wrong.. \[\int\limits_{?}^{?} y^4 +y^2-1/(y^3+y)] I went through and got y+ 1/y^3+y which leaves me at a dead end?

myininaya (myininaya):

I'm having trouble reading it... one sec

OpenStudy (anonymous):

sorry its \[\int\limits_{?}^{?} y^4+y^2-1/(y^3+y)\]

myininaya (myininaya):

\[\int\limits\limits_{?}^{?}\frac{ y^4 +y^2-1}{y^3+y} dy\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I do long divison to get y+ 1/y^3+y

myininaya (myininaya):